1) What is the volume of a 12.5 kg bar of gold?

2) The current price of gold is $40.18 for a gram or $1,249.65 per oz. What is the current price of the bar of gold in problem 1?

3) What is the volume of a 1.0 kg of silver?
Answer #1

1) Density off gold = d = 19.32 g/cm3 = 19320 kg/m3

Volume = Mass/Density = 12.5/19320 = 0.000647 m3

Hence the volume =   0.000647 m3

2) Price of 12.5 kg (12500 g gold) = 12500 x 40.18 = $502250

3) Density of silver = 10500 kg/m3

Volume = 1/10500 = 0.0000952 m3

Hence the volume of 1 kg of silver is   0.0000952 m3